### v2

# Spectroscopic readout of chiral photonic topology in a single-cavity spin-orbit-coupled Bose–Einstein condensate

Abstract
Topological photonic phases are typically identified through band reconstruction, steady-state transmission, or real-space imaging of edge modes. In this work, we present a framework for spectroscopic readout of chiral photonic topology in a single driven optical cavity containing a spin-orbit-coupled Bose-Einstein condensate. We demonstrate that the cavity transmission power spectral density provides a direct and measurable proxy for a momentum- and frequency-resolved photonic Chern marker, enabling topological characteristics to be inferred from spectral data without the need for bulk-band tomography. In the loss-dominated regime, where cavity decay exceeds atomic dissipation, the power spectral density exhibits Dirac-like gapped hybrid modes with a vanishing Chern marker, indicating a trivial phase. When the dissipation imbalance is reversed, a bright, gap-spanning spectral ridge emerges, co-localized with peaks in both the Chern marker and Berry curvature. The complex spectrum reveals parity-time symmetric coalescences and gain-loss bifurcations, marking exceptional points and enabling chiral, gap-traversing transport. By linking noise spectroscopy to geometric and non-Hermitian topology in a minimal cavity-QED architecture, this work provides a framework for spectroscopic detection of topological order in driven quantum systems. This approach offers a pathway to compact, tunable topological photonics across a broad range of light-matter platforms, providing a method for the study and control of topological phases in hybrid quantum systems.
Topological photonics has enabled unprecedented control of light by exploiting global geometric phases rather than local material properties Lu2014 ; Ozawa2019 . Analogues of quantum Hall and quantum spin Hall insulators have been realized in photonic crystals, coupled resonator arrays and metamaterials, revealing robust chiral edge modes that propagate without backscattering Hafezi2013 ; Rechtsman2013 ; Khanikaev2013 ; Barik2018 . More recently, attention has shifted to non-Hermitian and driven photonic systems, where balanced gain and loss, parity–time (PT) symmetry and exceptional points (EPs) give rise to topological phases without Hermitian counterparts ElGanainy2018 ; Ozdemir2019 ; Cerjan2019 ; Liu2025 . In these platforms, the real and imaginary parts of complex eigenvalues jointly determine topology, yet most experimental probes rely on steady-state transmission or band tomography and provide limited access to local signatures such as Berry curvature or real-space Chern markers Parto2021 ; Hu2021 ; Miri2019 . A method to extract topological information directly from cavity output fields, including fluctuation and noise spectra, is still lacking.
Ultracold atoms in optical cavities provide a natural interface between quantum optics and topological matter Ritsch2013 ; Mivehvar2021 . In particular, spin–orbit coupled (SOC) Bose–Einstein condensates (BECs) mimic solid-state topological phases, exhibiting Dirac dispersions, magnetic textures and Berry curvature hotspots Lin2011 ; Galitski2013 ; Zhai2015 ; Lian2017 . When placed inside a high-finesse cavity, such atoms hybridize with the cavity mode, allowing their internal pseudo-spin structure to be imprinted onto light Brennecke2007 ; Padhi2014 ; Deng2014 ; Ostermann2021 . This has enabled observations of cavity-induced magnetism, superradiant phase transitions and optomechanical backaction modified by SOC Hamner2014 ; Yasir2017 ; Landig2015 ; yasir2023 ; yasir2024 . Topological features have been inferred from mean-field cavity transmission, electromagnetically induced transparency (EIT) and Raman spectroscopy Clerk2010 ; Jia2023 ; Yasir2022 ; Qin2021 ; Peng2014 . However, these approaches primarily access average transmission or global Chern numbers in quasi-Hermitian regimes, and do not resolve how topology, dissipation and quantum fluctuations manifest in the power spectral density (PSD) of the cavity output. Moreover, the role of non-Hermitian band topology, Berry curvature and exceptional points in cavity-coupled SOC BECs remains largely unexplored Minganti2019 ; Lieu2018 ; Chen2022 .
In this article, we demonstrate that cavity transmission PSD offers a direct, experimentally accessible probe of non-Hermitian topology in a driven SOC BEC–cavity system. By computing the momentum- and frequency-resolved PSD of the output field, including quantum and thermal fluctuations, we construct a photonic Chern marker, a local analogue of the Chern number that characterizes topology without requiring integration over the full Brillouin zone Bianco2011 ; Huang2014 ; Marzari2012 . When cavity decay dominates atomic dissipation ($\kappa>\gamma$), the hybrid atom-photon spectrum exhibits gapped Dirac-like bands, and the Chern marker forms localized hotspots near avoided crossings. In contrast, when $\gamma>\kappa$, the system enters a $\mathcal{PT}$-symmetric gain-loss regime, where real and imaginary eigenvalues coalesce at exceptional points and then bifurcate to form ring-shaped gapless contours around the pseudo-spin manifolds ElGanainy2018 ; Cerjan2019 . The corresponding Chern marker reorganizes into annular, sign-alternating structures, tracing edge-like photonic modes. Additionally, we reconstruct the Berry curvature from the complex band structure inferred from the PSD and show its correspondence with the Chern marker distribution. Our work establishes cavity transmission noise as a powerful tool for observing Chern markers, Berry curvature, and exceptional-point physics in hybrid atom-photon systems, offering a pathway toward topological sensing and information processing with quantum gases in cavities.
Results
System Description
We analyze a spin–orbit–coupled Bose–Einstein condensate embedded in a single–mode optical resonator, motivated by the $^{87}\mathrm{Rb}$ implementations of Ref. Lin2011 . The atomic medium is a BEC with $N\simeq 1.8\times 10^{5}$ atoms confined inside a high–$Q$ Fabry–Pérot cavity of length $L\simeq 12.5\times 10^{-3}~\mathrm{m}$, see Fig. 1(a). The cavity axis is taken along $\hat{x}$ and supports a single mode of frequency $\omega_{c}\simeq 1.9\times 2\pi~\mathrm{GHz}$ with photon loss rate $\kappa$. A laser of power $P$ drives this mode along $\hat{x}$ with amplitude $|\eta|=\sqrt{P\kappa/(\hbar\omega_{E})}$, where the drive frequency is written as $\omega_{E}=\omega_{R}+\delta\omega_{R}$ to make explicit its relation to the Raman fields. The corresponding detuning between the pump and the bare cavity resonance is $\Delta_{c}=\omega_{E}-\omega_{c}$, and we operate in a near–dispersive regime with $\Delta_{c}\approx\kappa$, such that $\omega_{c}\simeq\omega_{R}+\delta\omega_{R}$ in the limit $\kappa\rightarrow 0$.
A homogeneous magnetic bias field of magnitude $B_{0}\simeq 10~\mathrm{G}$ is applied perpendicular to the cavity axis (in the yyzz plane), producing a Zeeman splitting $\hbar\omega_{z}$ between selected hyperfine levels and ensuring $\lvert\omega_{z}/\kappa\rvert\gg 1$. The pseudo–spin degree of freedom is formed by the two internal states $|\uparrow\rangle=|F=2,m_{F}=0\rangle$ and $|\downarrow\rangle=|F=2,m_{F}=-1\rangle$ within the $F=2$ manifold. SOC is realized using two counter–propagating Raman beams of frequencies $\omega_{R}$ and $\omega_{R}+\delta\omega_{R}$ and wavelength $\lambda=804.1~\mathrm{nm}$, propagating along $\hat{x}$ with wave vectors $\pm k_{x}\hat{x}$, where $k_{x}=2\pi/\lambda$. The single–photon recoil energy is $E_{x}=\hbar^{2}k_{x}^{2}/(2m_{a})$ for atomic mass $m_{a}$, and the two–photon Raman detuning is chosen as $\delta\simeq 1.6E_{x}$, in accordance with Ref. Lin2011 . In this geometry the Raman process ties the atomic spin to the center–of–mass motion along $\hat{x}$, generating an equal Rashba–Dresselhaus SOC in which the two spin components experience shifted dispersion minima and momentum–dependent spin mixing.
The single–particle dynamics of the pseudo–spin states are captured by the SOC Hamiltonian $\hat{\mathcal{H}}_{SOC}=\frac{\hbar^{2}\boldsymbol{k}^{2}}{2m_{a}}\sigma_{0}+\tilde{\alpha}k_{x}\sigma_{y}+\frac{\delta}{2}\sigma_{y}+\frac{\Omega_{z}}{2}\sigma_{z},$ where the quasimomentum is effectively one–dimensional, $\boldsymbol{k}=(k_{x},0,0)$, due to tight confinement. Here $\sigma_{0}$ is the $2\times 2$ identity, $\sigma_{x,y,z}$ are Pauli matrices acting in the $\{|\uparrow\rangle,|\downarrow\rangle\}$ basis, and the SOC strength is $\tilde{\alpha}=E_{x}/k_{x}=\hbar^{2}k_{x}/(2m_{a})$. The parameters $\delta$ and $\Omega_{z}$ are expressed as $\delta=-g\mu_{B}B_{z}$ and $\Omega_{z}=-g\mu_{B}B_{y}$, with $g$ the Landé gg–factor, $\mu_{B}$ the Bohr magneton, and $B_{y}$ and $B_{z}$ the components of the bias field along $\hat{y}$ and $\hat{z}$, respectively. In this language, $\delta$ encodes a tunable Zeeman field along $\hat{y}$, while $\Omega_{z}$ represents an effective Zeeman field along $\hat{z}$ controlled by the bias–field orientation.
Within the rotating–wave approximation, the many–body dynamics of the coupled BEC–cavity system are described by
$\hat{\mathcal{H}}$ $=$ $\int dx\,\boldsymbol{\hat{\psi}}^{{\dagger}}(x)\big[\hat{\mathcal{H}}_{0}+\mathcal{V}(x)\big]\boldsymbol{\hat{\psi}}(x)+\frac{1}{2}\int dx\sum_{\sigma,\sigma^{\prime}}\mathcal{U}_{\sigma\sigma^{\prime}}\,\hat{\psi}^{{\dagger}}_{\sigma}\hat{\psi}^{{\dagger}}_{\sigma^{\prime}}\hat{\psi}_{\sigma^{\prime}}\hat{\psi}_{\sigma}$ (1)
$+\hbar\Delta_{c}\,\hat{c}^{{\dagger}}\hat{c}-i\hbar\eta(\hat{c}-\hat{c}^{{\dagger}}),$
where $\boldsymbol{\hat{\psi}}=(\hat{\psi}_{\uparrow},\hat{\psi}_{\downarrow})^{T}$ denotes the two–component bosonic field operator, $\hat{c}$ ($\hat{c}^{{\dagger}}$) annihilates (creates) a cavity photon, and $\hat{\mathcal{H}}_{0}\equiv\hat{\mathcal{H}}_{SOC}$ is the single–particle Hamiltonian defined above.
The atoms are subject to a dispersive cavity-induced optical potential $\mathcal{V}(x)=\hbar U_{0}\,\hat{c}^{{\dagger}}\hat{c}\cos^{2}(k_{c}x)$, where $U_{0}=g_{0}^{2}/\Delta_{a}$ is the light shift per intracavity photon (with $g_{0}$ the single–photon Rabi frequency and $\Delta_{a}$ the atom–cavity detuning) and $k_{c}$ denotes the cavity wave vector. This potential originates from the standing-wave cavity mode and is dynamically controlled by the intracavity photon number $\hat{c}^{{\dagger}}\hat{c}$. In the collinear configuration considered here, the cavity field and the Raman beams propagate along the same axis but are spectrally and polarization selective: the control field populates the cavity mode, while the Raman fields implement the two-photon coupling between $|\uparrow\rangle$ and $|\downarrow\rangle$ with negligible cavity excitation. Any scalar AC Stark shift from the Raman beams is independent of $\hat{c}^{{\dagger}}\hat{c}$ and can be absorbed into the chemical potential (or treated as a weak static background), so $\mathcal{V}(x)$ provides the relevant dynamical lattice potential that enters our effective description Lin2011 ; Brennecke2007 ; haldane1988 ; Kane2005 ; Hasan2010 .
Contact interactions between atoms in spin components $\sigma,\sigma^{\prime}\in\{\uparrow,\downarrow\}$ are modeled by $\mathcal{U}_{\sigma\sigma^{\prime}}=4\pi\hbar^{2}a_{\sigma\sigma^{\prime}}/m_{a}$, where $a_{\sigma\sigma^{\prime}}$ denote the ss–wave scattering lengths. This form incorporates both density–density and spin–exchange channels within the SOC BEC coupled to the dynamical cavity field.
We assume $U_{\uparrow\uparrow}=U_{\downarrow\downarrow}=U$ and $U_{\uparrow\downarrow}=U_{\downarrow\uparrow}=\varepsilon U$. Using the plane-wave ansatz $\boldsymbol{\hat{\psi}}(x)=e^{i\boldsymbol{k}\cdot x}\boldsymbol{\hat{\varphi}}$, with $\boldsymbol{\hat{\varphi}}=(\hat{\varphi}_{\uparrow},\hat{\varphi}_{\downarrow})^{T}$ normalized as $|\hat{\varphi}_{\uparrow}|^{2}+|\hat{\varphi}_{\downarrow}|^{2}=N$, the dynamics reduce to the coupled quantum–Langevin equations,
$\dot{\hat{c}}$ $=$ $(i\Delta_{a}-ig_{a}\boldsymbol{\hat{\varphi}}^{{\dagger}}\boldsymbol{\hat{\varphi}}-\kappa)\hat{c}+\eta+\sqrt{2\kappa}\,a_{\mathrm{in}},$
$\dot{\boldsymbol{\hat{\varphi}}}$ $=$ $\Big[\frac{\hbar\boldsymbol{k}^{2}}{2m_{a}}\sigma_{0}+\tilde{\alpha}k_{x}\sigma_{y}+\frac{\delta}{2}\sigma_{y}+\frac{\Omega_{z}}{2}\sigma_{z}-\gamma+g_{a}\hat{c}^{{\dagger}}\hat{c}\Big]\boldsymbol{\hat{\varphi}}$ (2)
$+\frac{U}{2}\boldsymbol{\hat{\varphi}}^{{\dagger}}\boldsymbol{\hat{\varphi}}\,\boldsymbol{\hat{\varphi}}+\frac{\varepsilon U}{2}\hat{\varphi}_{\sigma}^{{\dagger}}\hat{\varphi}_{\sigma^{\prime}}\hat{\varphi}_{\sigma}+\sqrt{2\gamma}\,f_{a},$
where $a_{\mathrm{in}}$ and $f_{a}$ are the input noise operators of the cavity and atomic fields.
The collective density excitations of the condensate act as two effective atomic oscillators with frequency $\Omega=\hbar k^{2}/m_{\mathrm{bec}}$, driven by radiation pressure. Linearizing Eqs. (System Description) (see supply ) yields the effective optomechanical coupling $G=\sqrt{2}\,g_{a}|c_{s}|$ and detuning $\Delta=\Delta_{a}+g_{a}N$, where $g_{a}=(\omega_{c}/L)\sqrt{\hbar/(m_{\mathrm{bec}}\Omega)}$ and $|c_{s}|$ is the steady-state intracavity amplitude. The quantity $m_{\mathrm{bec}}=\hbar\omega_{c}^{2}/(L^{2}U_{0}^{2}\Omega)$ denotes the effective mass of the atomic mirror.
We extract topology directly from the measured cavity transmission spectrum. Using input–output theory on frequency domain solution of linearized quantum–Langevin equations, PSD of the output field, directly encodes the photonic Chern marker of the hybrid atom–cavity system Silveirinha2019 , see supplementary materials supply and Methods for details.
$S_{\mathrm{out}}(P,\omega)=\Xi(\omega)\,E_{T,\omega}\,\frac{C(k,\omega)}{2\pi},\Xi(\omega)=\frac{2\,\eta_{\mathrm{det}}\,\kappa_{\mathrm{ext}}}{|R(\omega)|^{2}},$ (3)
where $R(\omega)$ is the cavity response function and $E_{T,\omega}=\frac{\hbar\omega}{2}\coth\!\left(\frac{\hbar\omega}{2k_{B}T}\right)$. $C(k,\omega)$ is the local Chern marker carrying information bulk transport on edge modes, as illustrated in following findings.
Chiral-Photonic topology and Chern Marker
Figure 1: Influence of SOC and interspecies interactions on the transmission PSD. (a)–(c) Transmitted-field PSD $S_{\rm out}(k,\omega)$ for increasing SOC strength $\alpha=1.5\,\Omega$ (a), $2.5\,\Omega$ (b), and $3.5\,\Omega$ (c), at fixed Raman and dissipation parameters (as in the main text). Larger $\alpha$ enhances spin–momentum locking, splits an almost degenerate response into two SOC-hybridized polaritonic branches, and strengthens the $\Omega_{z}$-controlled gap while increasing the $k$-asymmetry. (d)–(f) PSD at fixed $\alpha$ for increasing interspecies interaction ratio $\epsilon\equiv U_{\uparrow\downarrow}/U$: $\epsilon=0$ (d), $\epsilon=1$ (e), and $\epsilon=2$ (f). Increasing $\epsilon$ renormalizes the atomic dispersion and dispersive light shift, shifting ridge frequencies and reducing gap contrast as the system approaches the SU(2)-symmetric point $\epsilon=1$. Across all panels, bright ridges trace poles of the linearized response, while linewidths reflect the net damping set by $\kappa$ and $\gamma$.
First, we consider the loss-dominated regime ($\kappa>\gamma$), where topology is trivial (Fig. 1). In Fig. 1(a), the pump interrogates the SOC-dressed pseudo-spin manifold. The transmitted field encodes the hybrid response in its power spectral density (PSD), $S_{\mathrm{out}}(P,\omega)$, where the ridges track the poles of the linear resolvent, and the linewidths are primarily determined by $\kappa$.
In Fig. 1(b), with the Raman (effective Zeeman) coupling switched off ($\Omega_{z}=0$), the SOC does not imprint a $k$-resolved band structure on the optical readout. Consequently, the transmission reduces to a conventional cavity spectrum: two resolved, $\kappa$-broadened sidebands (Stokes/anti-Stokes-like quasiparticle features). The corresponding Chern marker in Fig. 1(d) vanishes across $(k,\omega)$, confirming the absence of edge spectral flow.
Turning on $\Omega_{z}=3\Omega$, see Fig. 1(c), activates SOC hybridization and opens a Dirac-like gap. The upper and lower polaritonic bands separate around the former crossing, with a gap that scales with $\Omega_{z}$. Since photon loss exceeds atomic dissipation ($\kappa>\gamma$), no gap-spanning ridge emerges, and no mode connects the bands across $k$. The reconstructed Chern marker in Fig. 1(e) remains near zero across $(k,\omega)$, certifying a topologically trivial response.
Physically, $\Omega_{z}$ mixes the pseudo-spin states and opens a mass gap, while $\kappa>\gamma$ suppresses the non-Hermitian PT-symmetric feedback required to amplify an edge channel or create spectral winding/exceptional points. Thus, the cavity acts as a bulk spectrometer: with $\Omega_{z}=0$, it shows standard sidebands; with $\Omega_{z}\neq 0$, it resolves an SOC-induced gap. In both cases, the PSD-derived Chern marker quantitatively certifies trivial topology.
Figure 2: Emergence of gap-spanning edge states and PSD-derived topology in the gain-dominated regime ($\gamma>\kappa$). a, Power spectral density $S_{\mathrm{out}}(k,\omega)$ at Raman coupling $\Omega_{z}=3\,\Omega$ for $\gamma>\kappa$. A bright, continuous branch traverses the bulk gap, revealing an edge mode that connects the upper and lower polaritonic bands across wavevector $k$. b, Magnified view of the gap region in a. The inset arrows (blue/red) indicate opposite group velocities along the traversing branch, consistent with two-way chiral transport supported by non-Hermitian gain–loss imbalance. c, Chern-marker distribution reconstructed from the PSD in a. Two well-resolved positive peaks align with the gap-crossing branch and encode opposite transport directions of the chiral edge modes (signalled by the local spectral slope). The vanishing background away from the branch indicates that topology is concentrated at the gap-spanning trajectory. d, Chern marker as a function of input pump power $P$ for selected wavevectors $k_{x}$. The black curve ($k_{x}=0$) corresponds to the bulk region where no edge mode exists, giving a vanishing marker at all powers. The blue ($k_{x}=11\times 2\pi$) and red ($k_{x}=31\times 2\pi$) curves track the momenta of the first and second edge modes, respectively, both showing a pronounced increase of the marker with power as the chiral edge channels emerge and intensify with stronger light–matter coupling.
The principal result of this work is that the topology of a driven SOC–BEC in a single optical cavity can be read out spectroscopically from the transmission PSD, without spatially resolving edges. In the gain-dominated regime ($\gamma>\kappa$), Fig. 2(a) shows that $S_{\mathrm{out}}(k,\omega)$ develops a bright, gap-spanning branch that connects the upper and lower polaritonic bands across the wavevector axis. This branch is the optical signature of an edge channel stabilized by the non-Hermitian imbalance: atomic dissipation feeding back through the cavity provides an effective amplification pathway that compensates photon leakage and enables a mode to thread the Raman-induced bulk gap. The enlarged view in Fig. 2(b) highlights the chiral nature of this transport: the inset arrows (blue/red) indicate opposite signs of the local spectral slope $\partial\omega/\partial k$, i.e., counter-propagating group velocities along the traversing ridge. This is expected when non-Hermitian coupling lifts reciprocity constraints and allows bidirectional chiral flow in frequency–wavevector space.
From the same transmission data, we reconstruct a frequency– and wavevector–resolved Chern marker, as shown in Fig. 2(c). Two narrow, positive lobes are pinned to the gap-spanning trajectory, reflecting the concentration of topological weight at the edge branch. Although the local marker is positive in both lobes (by construction of the PSD-based estimator), the opposite transport directions are encoded in the sign of the spectral slope and in the Berry-curvature distribution surrounding each lobe. Away from the edge trajectory, the marker collapses to zero, indicating that the bulk bands remain topologically inert while the edge mode carries the nontrivial winding. This one-to-one correspondence between a gap-crossing PSD ridge and a localized Chern-marker response establishes a direct, quantitative bridge between optical spectra and topology in a minimal (single-cavity) platform.
Figure 2(d) illustrates the dependence of the Chern marker on the cavity input power $P$ for specific wavevector values $k_{x}$, selected to correspond to distinct regions of the spectrum. The black curve, evaluated at $k_{x}=0$, lies deep within the bulk band gap where no edge state exists; consequently, the Chern marker remains zero for all powers, confirming the absence of topological activity. The blue curve, taken at $k_{x}=11\times 2\pi$, corresponds to the momentum where the first edge mode emerges. Here, the Chern marker exhibits a sharp rise as the input power increases—signifying the onset of a topologically nontrivial regime once the optomechanical coupling $G\propto|c_{s}|$ becomes strong enough to overcome cavity losses and stabilize the chiral edge transport. The red curve, recorded at $k_{x}=31\times 2\pi$, traces the location of the second edge mode, which activates at slightly higher powers and yields a second distinct peak in the Chern marker. Together, these $k_{x}$-resolved curves reveal how the topological response can be tuned and selectively activated by controlling the pump power, mapping the successive appearance of edge channels directly onto an experimentally measurable photonic observable.
In the broader context of topological photonics, prior observations of edge transport typically relied on real-space imaging of waveguide arrays, photonic crystals, or ring-resonator lattices. In contrast, the present approach reads out edge physics from a bulk transmission spectrum in a single cavity, leveraging driven–dissipative (non-Hermitian) physics to generate a gap-spanning spectral branch and using a PSD-based Chern marker to certify its topology. This combination—gap-traversing edge ridges in $S_{\mathrm{out}}(k,\omega)$, chiral flow evidenced by opposite spectral slopes, and a co-localized Chern-marker signal with power-tunable strength—constitutes the core novelty of our protocol and establishes PSD spectroscopy as a compact, experimentally accessible route to topological diagnostics in hybrid light–matter systems.
Berry-curvature and non-Hermitian topology
Figure 3: Berry-curvature cartography and non-Hermitian band topology. a, Berry-curvature density $\Omega_{B}(k,\omega)$ reconstructed from the PSD for $\Omega_{z}=3\,\Omega$, shown as a 3D surface with a base-plane density projection. Curvature concentrates along the gap-edge trajectories identified in transmission. b, Same as a for $\Omega_{z}=5\,\Omega$, revealing enhanced, more sharply localized curvature lobes and a redistribution of weight as the Raman-induced mass increases. c, Real parts of the eigenvalues $\mathrm{Re}\,\lambda(k)$ at $\Omega_{z}=3\,\Omega$, displaying coalescence points characteristic of $\mathcal{PT}$-symmetric band crossings. d, Imaginary parts $\mathrm{Im}\,\lambda(k)$ at the same parameters, showing gain–loss bifurcation at the coalescence points in c, thus identifying the exceptional points (EPs) that delimit the transition between unbroken and broken $\mathcal{PT}$ phases.
The Berry-curvature maps in Fig. 3(a,b) provide a momentum–frequency–resolved view of the geometric response underlying the PSD features reported in Figs. 12. For $\Omega_{z}=3,\Omega$, Fig. 3(a) shows that $\Omega_{B}(k,\omega)$ is narrowly concentrated along the gap-edge trajectories, forming ridge-like lobes in the 3D surface and in the base-plane density projection. Increasing the Raman coupling to $\Omega_{z}=5,\Omega$ [Fig. 3(b)] sharpens these lobes and shifts the curvature weight outward in $(k,\omega)$, consistent with an enlarged Dirac mass and steeper dispersions. The curvature remains strongly localized where the PSD indicates maximal spectral flow, providing a geometric counterpart to the Chern-marker peaks. This redistribution of the Berry curvature reveals how the geometric phase accumulates in momentum space and reflects the topological features of the system, with stronger curvature found at regions of enhanced spectral flow, highlighting the critical role of the curvature in non-Hermitian systems.
Figures 3(c,d) link this geometric structure to non-Hermitian band topology. At $\Omega_{z}=3,\Omega$, the real parts of the eigenvalues $\mathrm{Re},\lambda(k)$, Fig. 3(c), exhibit clear coalescences, while the imaginary parts $\mathrm{Im},\lambda(k)$, Fig. 3(d), bifurcate precisely at those points—the hallmark of exceptional points (EPs) in a $\mathcal{PT}$-symmetric spectrum. These EPs demarcate the transition between the unbroken phase (equal linewidths, distinct frequencies) and the broken phase (frequency coalescence with asymmetric gain–loss), and they occur proximate to the high-curvature ridges. Notably, the coalescence points in $\mathrm{Re},\lambda(k)$ and the gain–loss bifurcations in $\mathrm{Im},\lambda(k)$ occur at the same $(k,\omega)$ loci that host the Berry-curvature ridges and the PSD gap-traversing branch. This one-to-one co-location of EPs, curvature hotspots, and edge spectral flow confirms that the emergent chiral transport in the $\gamma>\kappa$ regime is governed by non-Hermitian criticality. This joint Berry–EP spectroscopy in a single-mode cavity constitutes a compact and experimentally accessible route to diagnosing geometric and non-Hermitian topology in driven light–matter systems.
Edge modes with asymmetric phase transition of SOC-BEC
Figure 4: Phase–detuning control of edge transport and Chern-marker localization. Upper row: a–c, Transmission PSD $S_{\mathrm{out}}(k_{x},\omega)$ for Raman detuning $\delta/\Omega=-1,\,0,\,+1$, respectively, at fixed $\Omega_{z}=5\,\Omega$ in the gain-dominated regime ($\gamma>\kappa$). A finite detuning breaks the $k_{x}\!\to\!-k_{x}$ symmetry of the SOC–BEC phase and re-positions the gap-spanning edge branch within the bulk gap: for $\delta=-\Omega$ it is biased to negative $k_{x}$ (a), for $\delta=0$ it is centered (b), and for $\delta=+\Omega$ it shifts to positive $k_{x}$ (c). Lower row: d–f, Corresponding Chern-marker profiles along $k_{x}$ (frequency-integrated around the gap) showing a rigid displacement of the topological weight that tracks the edge-branch drift: left-shift for $\delta=-\Omega$ (d), symmetric/centered for $\delta=0$ (e), and right-shift for $\delta=+\Omega$ (f). The marker magnitude remains comparable, indicating that detuning redistributes rather than quenches topology, enabling phase-bias control of chiral transport.
The data in Fig. 4 establish Raman detuning $\delta$ as a precise phase-bias knob that repositions chiral transport in momentum space without quenching topology. In the upper row, $S_{\mathrm{out}}(k_{x},\omega)$ shows that a nonzero $\delta$ breaks the $k_{x}\to-k_{x}$ inversion of the SOC manifold and shifts the gap-spanning edge ridge: for $\delta=-\Omega$, as shown in Fig. 4(a), the traversing branch is biased to negative $k_{x}$; for $\delta=0$, Fig. 4(b) shows the ridge centered; and for $\delta=+\Omega$, Fig. 4(c) shows the ridge shifted to positive $k_{x}$ Lin2011 . This drift is consistent with the SOC Hamiltonian used here, where $\delta$ acts as an effective Zeeman term that tilts the Dirac mass and displaces the spectral flow in $(k_{x},\omega)$. Under gain-dominated conditions ($\gamma>\kappa$), the non-Hermitian imbalance then selects the shifted branch for amplification, yielding a bright gap-spanning ridge.
The lower row quantifies the same effect geometrically. Frequency-integrated Chern-marker profiles along $k_{x}$ [Figs. 4(d–f)] exhibit a rigid lateral translation that tracks the PSD ridge: the peak of the marker migrates left for $\delta=-\Omega$, remains centered for $\delta=0$, and moves right for $\delta=+\Omega$. Importantly, the marker amplitude remains nearly unchanged, indicating that detuning redistributes topological weight in momentum space rather than diminishing it. To leading order, this behavior is expected from a curvature-conservation picture: the Berry curvature condenses near the gap-crossing trajectory, and $\delta$ shifts where the non-Hermitian spectral winding is maximal, but the integrated weight (set by the edge-channel topology) is preserved.
Operationally, this realizes momentum-space routing of chiral transport in a single-mode cavity: by tuning $\delta$, one steers the $k_{x}$ at which the edge channel carries peak flow, providing a compact alternative to spatially engineered non-reciprocity in extended photonic lattices and $\mathcal{PT}$-symmetric platforms Baumann2010 ; harari2018 ; bandres2018 ; Bergholtz2021 .
From an experimental standpoint, the joint PSD/marker readout offers a direct calibration loop: the displacement of the PSD ridge with $\delta$ gives a spectroscopic handle on the effective Zeeman bias, while the co-moving Chern-marker peak certifies that the observed transport is genuinely topological rather than a dispersive artifact. The combination of phase-bias control (via $\delta$) and gain–loss selectivity (via $\gamma/\kappa$) thus enables deterministic placement and tuning of chiral edge channels in momentum space, leveraging non-Hermitian band topology in a minimal cavity-QED architecture.
Discussion
We have demonstrated that cavity transmission spectroscopy offers a quantitative and experimentally accessible method for diagnosing band topology in a driven SOC–BEC, using a single cavity. By extracting a frequency- and wavevector-resolved Chern marker from the transmitted PSD, a standard optical observable is transformed into a topological probe. Gap-traversing ridges in $S_{\mathrm{out}}(k,\omega)$ co-localize with Chern-marker peaks when edge transport is present, while purely bulk spectra yield a vanishing marker, providing a direct optical signature of topological edge modes.
Three key findings underpin this conclusion. (i) Dissipation-driven topology: In the loss-dominated regime ($\kappa>\gamma$), the Raman field opens a Dirac-like gap but no edge mode is stabilized, and the Chern marker remains zero. In the gain-dominated regime ($\gamma>\kappa$), a bright, gap-spanning ridge appears, with opposite spectral slopes (chiral flows) and a Chern marker concentrated along the trajectory of the edge mode. Power sweeps at selected $k_{x}$ values reveal the sequential activation of edge channels. (ii) Geometric and non-Hermitian certification: The Berry curvature reconstructed from the PSD condenses along gap edges and intensifies with increased Raman coupling, while real and imaginary eigenvalue pairs reveal $\mathcal{PT}$-symmetric exceptional points that delineate the transition between unbroken and broken phases. (iii) Phase-bias control: Raman detuning shifts the edge ridge and Chern marker across $k_{x}$ without reducing their amplitudes, enabling momentum-space routing of chiral transport within a minimal cavity-QED architecture.
These results establish cavity transmission PSD as a robust, quantitative probe of band topology in a driven SOC–BEC. The frequency- and wavevector-resolved Chern marker, extracted directly from $S_{\mathrm{out}}(P,\omega)$, co-localizes with gap-traversing ridges only in the gain-dominated regime ($\gamma>\kappa$), certifying the presence of edge transport. In the loss-dominated regime ($\kappa>\gamma$), the Chern marker vanishes, confirming the absence of edge states. Berry-curvature maps provide additional geometric insight, while $\mathcal{PT}$-symmetric eigenvalue coalescences identify exceptional points that govern the onset of chiral, gap-spanning transport.
The novelty of this work lies in using standard optical tools (PSD and Chern marker) to diagnose non-Hermitian topological features in a single-cavity system. This approach avoids the need for real-space imaging and instead relies on bulk transmission spectra to identify edge states and topological transport. The ability to tune and route chiral edge transport in momentum space via Raman detuning provides a compact and experimentally accessible method for controlling topological modes in driven light–matter systems.
In broader terms, this work fills a significant gap by demonstrating how non-Hermitian physics can be used to control and quantify topological transport in minimal cavity architectures, expanding the possibilities for future quantum photonic devices. Furthermore, this technique offers a pathway to studying geometric responses and non-Hermitian criticality in complex photonic systems with fewer experimental components compared to conventional lattice-based platforms. By extending this approach to multimode or Floquet cavities, we foresee the potential for developing spectroscopic topological photonics, where topological phases and edge channels can be engineered, read out, and tuned with standard optical methods, providing a powerful tool for quantum sensing, robust transport, and novel quantum optics applications.
Acknowledgements.
K.A.Y. acknowledges the support of Research Fund for International Young Scientists by NSFC under grant No. KYZ04Y22050, Zhejiang Normal University research funding under grant No. ZC304021914 and Zhejiang province postdoctoral research project under grant number ZC304021952.
Author contributions
The research was initiated and led by K.A.Y.; K.A.Y. and G.X.L. analyzed the results; K.A.Y wrote the paper under guidance G.X.L.
Code availability
All results illustrated in this word are generated from the known analytical expressions given in the article and its Supplemental Information. The detailed code is available from the corresponding author upon reasonable request.
Competing interests
The authors declare no competing interests.